OpenStudy (anonymous):

Factor the expression into a product of binomials. 4x^2-32xy+64y^2

OpenStudy (callisto):

4x^2-32xy+64y^2 = 4( x^2 - 8xy + 16y^2) = 4 ( x-4y)^2